What is the wavelength of light emitted by the transition of a hydrogen electron from n=3 to n=1 ?
a) 137 nm
b) 1.03 * 10^-7
c) 103 nm

A typed answer is prefered.
Answer #1

We know that the frequency ( ѵ) of the emitted light is given as-

phplbQP96.png

where R is the Rydberg constant, where R= 109737.316 cm-1  and n1 = 1 ; n2 = 3.

Putting the values and calculating we get  ѵ= 97544.280 cm-1

so wavelength ( λ ) = (1/ ѵ) cm = 102.5 x 10-7 cm = 102.5 nm which is closer to 103 nm

The answer will be c) 103 nm